The life span of one species of a maple tree is normally distributed with mean 20 years and standard deviation 2.3 years. If one tree is choosen at random, what is the probability it will live 21 years or longer? What is the probability that the average lifetime of 9 such trees will be 21 years or more?
Answer by stanbon(26297)

The life span of one species of a maple tree is normally distributed with mean 20 years and standard deviation 2.3 years. If one tree is choosen at random, what is the probability it will live 21 years or longer?
---
z(21) = (21-20)/2.3 = 0.4348
P(x >= 21) = P(z >= 0.4348) = 0.3319
----------------------------------------------

What is the probability that the average lifetime of 9 such trees will be 21 years or more?
---
z(21) = (21-20)/[2.3/Sqrt(9)] = 1/[2.3/3] = 3/2.3 = 1.305
P(x-bar >= 21) = P(z >= 1.305) = 0.09606
=============================================
Cheers,
Stan H.